OpenStudy (b77w):

How do I simplify this? ((2x-2y)/(3x+3y))*((6x+6y)/(x^2-y^2))

OpenStudy (campbell_st):

take out common factors and factor x^2 - y^2 to start

OpenStudy (campbell_st):

so it can be written as \[\frac{2(x -y)}{3(x + y)} \times \frac{6(x + y)}{(x -y)(x+y)}\] now cancel common factors and then multiply whats left